# FreeChunker: A Cross-Granularity Chunking Framework

### v1

@@ 5.2 Training Dynamics and Evaluation
The lower panels of Figure 3 illustrate the cosine loss distributions on the independent test set. We employ a threshold of 0.06 to distinguish well-fitted samples from under-fitted ones, enabling coarse-grained assessment of chunk embedding accuracy. For jina-embeddings-v2-small-en, 100% of test samples fall below this threshold, with loss distribution highly concentrated in the 0.01–0.02 range, reflecting excellent generalization capability and minimal approximation error. In contrast, nomic-embed-text-v1.5 (137M parameters) exhibits a more dispersed distribution: 96.7% of samples remain below the 0.06 threshold, yet long-tail high-loss samples emerge, indicating increased representation noise or model uncertainty. BGE-M3 (570M parameters) shows further distribution broadening, consistent with the training loss patterns observed.
@@ 6.1 Experimental Setup
Comprehensive experiments are conducted on LongBench V2 (Bai et al., 2024), a challenging benchmark specifically designed for long-context understanding tasks. This dataset is suitable for evaluating chunking methods as it contains documents with extensive contexts and covers diverse task types with varying difficulty levels, thereby providing a realistic testbed for assessing the effectiveness of different chunking strategies in practical RAG scenarios.
@@
For text generation, Qwen3-8B (Team, 2025) is employed as the backbone language model, deployed via vLLM for efficient inference. To guarantee fair comparison and reproducibility, greedy decoding with temperature set to 0 is employed by Qwen3-8B.
@@
FreeChunker is compared against several representative chunking approaches. (1) Traditional Chunking represents the traditional approach that first splits text by periods to preserve semantic integrity, then iteratively accumulates sentences until exceeding predefined token limits (256 tokens); (2) Semantic Chunking employs SemanticChunker (LangChain, 2024) to perform embedding similarity-based chunking using sentence embeddings to identify natural breakpoints in the text. The breakpoint threshold type set to "percentile" and breakpoint threshold amount set to 50.0; (3) Meta-Chunking (Zhao et al., 2024) employs perplexity-based semantic boundary detection with two sub-methods: PPL Chunking and Margin Sampling, utilizing Qwen/Qwen2.5-1.5B-Instruct model (Yang et al., 2024) for text chunking. PPL Chunking adopts the threshold of 0.5 as recommended in the original paper, while Margin Sampling uses default settings; (4) LumberChunker (Duarte et al., 2024) provides LLM-guided chunk boundary prediction; however, due to excessive processing time for long texts, it is excluded from comparisons involving lengthy documents.
@@ 6.2 Main Results
Table 1 presents the results across three embedding models on LongBench V2. FreeChunker demonstrates superior performance in complex query environments while achieving significant computational efficiency gains compared to traditional approaches.
Regarding computational efficiency, FreeChunker maintains processing time comparable to traditional chunking methods while delivering superior retrieval performance. In contrast, existing semantic chunking approaches require significantly longer processing time without achieving comparable performance gains. This combination of enhanced effectiveness and computational efficiency validates the core contributions of the cross-granularity paradigm in practical deployment scenarios.
@@ 7 Conclusion
This paper introduces FreeChunker, a novel cross-granularity chunking framework that fundamentally transforms the traditional text chunking paradigm. By shifting from static boundary identification to flexible semantic combination retrieval, the proposed method significantly reduces computational overhead associated with semantic analysis for boundary detection while providing enhanced adaptability to diverse query requirements. Unlike conventional static chunking approaches, this flexible semantic combination strategy accommodates varying and unpredictable chunking demands across different query intents.
Beyond the adjacent sentence combination approach explored in our experiments, the framework’s chunk pattern control mechanism $\mathbf{P}_{g,s}[i,j]$ enables non-contiguous sentence linking capabilities. This opens promising research directions for incorporating predefined network structures within document collections to construct sophisticated non-adjacent sentence connections. Such extensions represent compelling avenues for future investigation, potentially enabling more complex semantic relationship modeling across document structures.
The experimental results demonstrate that FreeChunker achieves competitive retrieval performance while maintaining computational efficiency comparable to baseline methods. The framework’s ability to provide multiple granularity levels simultaneously addresses fundamental limitations of existing single-granularity approaches, offering enhanced flexibility for diverse information retrieval scenarios in practical RAG applications.

### v2

@@ 4.2 Training Dynamics and Evaluation
The lower panels of Figure 3 illustrate the cosine loss distributions on the independent test set. A threshold of 0.06 is employed to distinguish well-fitted samples from under-fitted ones, enabling coarse-grained assessment of chunk embedding accuracy. For jina-embeddings-v2-small-en, 100% of test samples fall below this threshold, with loss distribution highly concentrated in the 0.01–0.02 range, reflecting excellent generalization capability and minimal approximation error. In contrast, nomic-embed-text-v1.5 (137M parameters) exhibits a more dispersed distribution: 96.7% of samples remain below the 0.06 threshold, yet long-tail high-loss samples emerge, indicating increased representation noise or model uncertainty. BGE-M3 (570M parameters) shows further distribution broadening, consistent with the training loss patterns observed.
5 Theoretical Analysis
While Section 4 demonstrates that the cross-granularity encoder achieves a low fitting loss (i.e., high cosine similarity $\rho\approx 1$ between true and approximate embeddings), it remains crucial to understand how this approximation error propagates to the downstream retrieval task.
5.1 Setup and Constraints
Let $\vec{e}_{i,j}=\mathcal{M}([s_{i},\dots,s_{j}])$ denote the true chunk embeddings of the concatenated chunk $s_{i},\dots,s_{j}]$, and let $\vec{v}_{i,j}$ be the approximate chunk embeddings produced by the cross-granularity encoder. Let $\vec{q}$ be a query embedding. Throughout this section all vectors are assumed to be unit-norm.
The substitution loss is studied as follows:
Two parameters characterize the geometric configuration:
$\rho$ $\;=\;\cos(\vec{e}_{i,j},\vec{v}_{i,j}),$ (6)
$s$ $\;=\;\cos(\vec{q},\vec{e}_{i,j}),$
5.2 Upper Bound Analysis
To rigorously quantify the retrieval quality in high-dimensional embedding spaces (e.g., $d\geq 768$), a probabilistic bound is established by leveraging the Concentration of Measure phenomenon. This derivation proves that the probability of a random query aligning with a specific approximation error direction is exponentially small.
Step 1: Orthogonal Decomposition
A coordinate system is established aligned with the true chunk embedding $\vec{e}_{i,j}$. Any unit vector $\vec{x}$ can be decomposed into a component parallel to $\vec{e}_{i,j}$ and a component in the orthogonal complement $\vec{e}_{i,j}^{\perp}$: $\vec{x}=(\vec{x}\cdot\vec{e}_{i,j})\vec{e}_{i,j}+\vec{x}_{\perp}$.
Given $\vec{v}_{i,j}\cdot\vec{e}_{i,j}=\rho$, the magnitude of the orthogonal component is $\sqrt{1-\rho^{2}}$. It can be written as:
$\vec{v}_{i,j}=\rho\vec{e}_{i,j}+\sqrt{1-\rho^{2}}\,\mathbf{u},$ (7)
where $\mathbf{u}\in\mathbb{R}^{d}$ is a unit vector in the orthogonal subspace ($\mathbf{u}\perp\vec{e}_{i,j}$). Intuitively, $\mathbf{u}$ represents the direction of the approximation error.
Given $\vec{q}\cdot\vec{e}_{i,j}=s$, the query is similarly decomposed:
$\vec{q}=s\vec{e}_{i,j}+\sqrt{1-s^{2}}\,\mathbf{w},$ (8)
where $\mathbf{w}\in\mathbb{R}^{d}$ is a unit vector in the orthogonal subspace ($\mathbf{w}\perp\vec{e}_{i,j}$), determined solely by the query.
Step 2: Algebraic Expansion of Loss
Substituting the decomposed forms into the dot product $\vec{q}\cdot\vec{v}_{i,j}$:
$\vec{q}\cdot\vec{v}_{i,j}$ $=(s\vec{e}_{i,j}+\sqrt{1-s^{2}}\mathbf{w})\cdot(\rho\vec{e}_{i,j}+\sqrt{1-\rho^{2}}\mathbf{u})$
$=s\rho(\vec{e}_{i,j}\cdot\vec{e}_{i,j})+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}(\mathbf{w}\cdot\mathbf{u})$
$=s\rho+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}(\mathbf{w}\cdot\mathbf{u}).$ (9)
Note that cross-terms vanish due to orthogonality. Substituting this back into the loss expression $\varepsilon=|s-\vec{q}\cdot\vec{v}_{i,j}|$:
$\varepsilon$ $=\left|s-\left(s\rho+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}(\mathbf{w}\cdot\mathbf{u})\right)\right|$
$=\left|s(1-\rho)-\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}(\mathbf{w}\cdot\mathbf{u})\right|.$ (10)
Using the triangle inequality $|a-b|\leq|a|+|b|$, the structural upper bound is obtained:
$\varepsilon\leq\underbrace{s(1-\rho)}_{\text{Radial Error}}+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}\underbrace{|\mathbf{w}\cdot\mathbf{u}|}_{\text{Tangential Noise}}.$ (11)
Step 3: Probabilistic Bound via Concentration
The term $|\mathbf{w}\cdot\mathbf{u}|$ represents the projection of the error direction $\mathbf{u}$ onto the query’s specific orthogonal direction $\mathbf{w}$. Both vectors reside in a (d1)(d-1)-dimensional subspace.
Assumption (Quasi-Random Error). For an independent query, the direction of the approximation error $\mathbf{u}$ is uniformly distributed on the sphere $S^{d-2}$ relative to $\mathbf{w}$.
While $\mathbf{u}$ arises from aggregation encoding, it is generally decoupled from the semantic intent of user queries. In high-dimensional manifolds ($d\gg 100$), the probability of an independent query aligning with a specific structural artifact vanishes, rendering them effectively orthogonal.
To quantify this, Levy’s Lemma (Milman and Schechtman, 1986; Ledoux, 2001) is applied. For a random vector $\mathbf{u}$ uniformly distributed on the sphere and a fixed vector $\mathbf{w}$, the probability that their projection magnitude $|\mathbf{w}\cdot\mathbf{u}|$ exceeds a threshold $t$ decays exponentially:
$P(|\mathbf{w}\cdot\mathbf{u}|\geq t)\leq 2\exp\left(-\frac{(d-1)t^{2}}{2}\right).$ (12)
Let $\delta\in(0,1)$ be a small failure probability. By equating the tail probability bound to $\delta$, the threshold $t$ is derived as follows:
$\begin{split}2\exp\left(-\frac{(d-1)t^{2}}{2}\right)&=\delta\\ \Longrightarrow\quad t&=\sqrt{\frac{2\ln(2/\delta)}{d-1}}.\end{split}$ (13)
This derivation implies that the projection error $|\mathbf{w}\cdot\mathbf{u}|$ is confined within this threshold $t$ with a confidence level of at least $1-\delta$.
Theorem (Probabilistic Substitution Bound). With probability at least $1-\delta$, the substitution loss is bounded by:
$\varepsilon\leq s(1-\rho)+\sqrt{1-s^{2}}\sqrt{1-\rho^{2}}\cdot\sqrt{\frac{2\ln(2/\delta)}{d-1}}.$ (14)
Implications.
This result highlights the “blessing of dimensionality”: as $d$ increases (e.g., $d\to\infty$), the tangential noise term $O(1/\sqrt{d})$ vanishes. The substitution loss is thus dominated by the radial term $s(1-\rho)$. Given that the query similarity $s$ is typically less than 1 ($s<1$) and the fitting process ensures $\rho$ is very close to 1 ($\rho\approx 1$), the term $s(1-\rho)$ becomes negligible. This confirms that the retrieval quality of the fitted vectors is theoretically guaranteed, verifying the robustness of FreeChunker in high-dimensional embedding spaces.
Embeddings Traditional SemanticChunker PPL Chunking Margin Sampling LumberChunker FreeChunker
Top-5 Top-10 Top-5 Top-10 Top-5 Top-10 Top-5 Top-10 Top-5 Top-10 Top-5 Top-10
I. Single-Document QA
jina-small-en 30.86 35.43 33.71 33.71 34.10 32.57 28.57 34.29 32.00 35.43 35.43 32.57
nomic-text-v1.5 29.90 29.71 29.71 31.43 32.19 33.71 32.57 31.81 29.52 29.14 35.62 32.57
BGE-M3 33.33 37.71 32.57 31.43 33.33 34.29 32.57 33.14 35.81 36.38 36.57 34.29
II. Multi-Document QA
jina-small-en 24.80 28.00 28.80 28.80 28.00 28.27 29.60 27.20 25.87 28.80 26.40 26.40
nomic-text-v1.5 26.40 26.40 22.40 24.00 25.07 27.20 28.00 26.40 25.60 28.80 24.00 28.00
BGE-M3 27.20 28.80 25.87 31.20 27.20 28.00 27.20 29.60 28.80 29.60 28.00 29.60
III. Code Repository Understanding
jina-small-en 44.00 51.33 36.00 44.00 38.00 48.00 42.00 48.00 44.00 52.67 46.00 44.00
nomic-text-v1.5 42.00 46.00 38.00 36.00 35.33 38.00 44.00 42.00 32.00 36.00 34.00 42.00
BGE-M3 42.00 44.00 44.00 36.67 48.00 46.00 40.00 36.00 40.67 38.00 42.00 48.00
IV. Long In-context Learning
jina-small-en 27.57 30.86 29.63 29.63 28.40 28.40 25.93 20.99 32.10 29.63 32.10 28.40
nomic-text-v1.5 28.40 28.40 28.40 29.63 26.75 19.75 26.75 22.22 22.63 29.63 20.99 24.69
BGE-M3 26.75 25.10 34.57 30.86 27.57 28.40 32.10 30.45 28.40 25.93 27.16 25.93
V. Long-dialogue History Understanding
jina-small-en 26.50 33.33 25.64 17.95 23.08 16.24 28.21 25.64 30.77 23.08 28.21 23.08
nomic-text-v1.5 35.04 25.64 30.77 20.51 20.51 17.95 23.08 33.33 23.08 25.64 38.46 38.46
BGE-M3 23.08 23.08 33.33 29.06 14.53 30.77 25.64 37.61 25.64 20.51 33.33 33.33
VI. Long Structured Data Understanding
jina-small-en 15.15 24.24 25.25 29.29 33.33 33.33 30.30 36.36 26.26 21.21 42.42 45.45
nomic-text-v1.5 35.35 30.30 27.27 21.21 33.33 37.37 39.39 39.39 12.12 12.12 36.36 33.33
BGE-M3 30.30 38.38 23.23 27.27 33.33 33.33 33.33 30.30 21.21 16.16 45.45 42.42
VII. Overall
jina-small-en 28.76 33.53 30.88 31.34 31.15 31.15 29.82 31.21 31.21 32.67 33.60 31.61
nomic-text-v1.5 30.75 30.02 28.43 28.23 29.03 29.29 31.34 30.55 26.04 28.43 30.48 31.61
BGE-M3 30.62 33.00 31.81 31.35 30.88 32.61 31.41 32.27 31.61 30.62 33.80 33.60
Average 30.04 32.18 30.37 30.31 30.35 31.01 30.86 31.35 29.62 30.57 32.63 32.27
Table 1: Average accuracy (%) of different chunking methods on LongBench V2. "Overall" denotes the average accuracy across all questions. Standard deviations (SDs) across 3 runs are minimal: 78.7% of accuracy SDs are 0.00%. Specifically, FreeChunker show all accuracy SDs << 0.50%, while others range up to 3.50%. The best and second-best results for FreeChunker and the Average score are highlighted in bold and underlined, respectively.
@@ 6.1 Experimental Setup
Comprehensive experiments are conducted on LongBench V2 Bai et al. (2024), a challenging benchmark specifically designed for long-context understanding tasks. This dataset is suitable for evaluating chunking methods as it contains documents with extensive contexts and covers diverse task types with varying difficulty levels, thereby providing a realistic testbed for assessing the effectiveness of different chunking strategies in practical RAG scenarios.
@@
For text generation, Qwen3-8B Team (2025) is employed as the backbone language model, deployed via vLLM Kwon et al. (2023) for efficient inference. To guarantee fair comparison and reproducibility, greedy decoding with temperature set to 0 is employed by Qwen3-8B.
Figure 4: Average processing time per instance across baselines on LongBench V2. The processing time comprises the sum of chunking and encoding durations. FreeChunker (highlighted in orange) achieves comparable speed to Traditional Chunking and significantly outperforms other methods.
@@
FreeChunker is compared against several representative chunking approaches. (1) Traditional Chunking represents the traditional approach that first splits text by periods to preserve semantic integrity, then iteratively accumulates sentences until exceeding predefined token limits (256 tokens); (2) Semantic Chunking employs SemanticChunker LangChain (2024) to perform embedding similarity-based chunking using sentence embeddings to identify natural breakpoints in the text. The breakpoint threshold type is set to "percentile" and the breakpoint threshold amount is set to 50.0; (3) Meta-Chunking Zhao et al. (2024) employs perplexity-based semantic boundary detection with two sub-methods: PPL Chunking and Margin Sampling, utilizing Qwen/Qwen2.5-1.5B-Instruct model Yang et al. (2024) for text chunking. PPL Chunking adopts the threshold of 0.5 as recommended in the original paper, while Margin Sampling uses default settings; (4) LumberChunker Duarte et al. (2024) provides LLM-guided chunk boundary prediction, utilizing the same configuration as the Generative Model.
FreeChunker Setting
The Sentenizer module adopts Traditional Chunking as the foundational sentence splitting method, with the same setting. For the granularity configuration, the simplest combination of three granularities is directly preset: (1, 2, 4). This setup ensures efficient processing while providing multi-scale context coverage.
@@ 6.2 Main Results
Overall Performance
Table 1 presents the comparative results on LongBench V2 across three embedding models. FreeChunker achieves the best average performance among all methods.
Ablation Study
Since FreeChunker utilizes Traditional Chunking as the underlying chunking method for its Sentenizer to segment basic sentence units, and subsequently applies the Cross-Granularity Encoder to obtain cross-granularity embeddings, the performance gap between them can be regarded as an ablation study for the proposed cross-granularity framework. As shown in Table 1, FreeChunker consistently outperforms the Traditional baseline, which verifies the superposition effect of multi-granularity embeddings and the effectiveness of the cross-granularity framework.
Time Efficiency
Figure 4 illustrates the time cost distribution across different tasks. Compared to Traditional Chunking, FreeChunker merely adds the forward propagation encoding time of the Cross-Granularity Encoder, resulting in negligible extra overhead. In contrast, LLM-based methods (e.g., LumberChunker) consume hundreds of seconds. This achieves a speedup of up to $30\times$ compared to complex semantic chunkers, demonstrating that FreeChunker effectively circumvents the efficiency bottleneck of semantic-aware chunking without compromising retrieval quality.
@@ 7 Conclusion
This paper introduces FreeChunker, a novel cross-granularity chunking framework that fundamentally transforms the traditional text chunking paradigm. By shifting from “static boundary identification and re-encoding” to “flexible semantic combination of sentence encodings,” this framework significantly reduces the computational overhead of semantic analysis required for boundary detection, while enhancing adaptability to diverse query requirements. Experimental results obtained using the simplest preset chunk patterns demonstrate significant advantages over existing methods, thereby validating the potential of this framework.
Beyond the preset chunk patterns explored in the current study, the framework’s chunk pattern control mechanism, $\mathbf{P}_{g,s}[i,j]$, provides non-contiguous sentence linking capabilities. By supporting sophisticated combination methods based on sentence encodings, the framework enables the construction of complex non-adjacent sentence connections and patterns tailored to specific scenarios and tasks. These capabilities offer promising prospects for future research into adaptive and structural document modeling.
Limitations
Although FreeChunker requires increased GPU memory (VRAM) for encoding due to the introduction of the Cross-Granularity Encoder, this overhead is manageable. Since the encoding operates at the sentence level, the input sequence length for the attention mechanism is significantly shorter compared to token-level approaches, resulting in efficient computation.
Ethics Statement
This work does not involve human subjects, animal experiments, or the collection of private or sensitive data. All datasets utilized in this research are publicly available and used in accordance with their respective licenses.
During the preparation of this manuscript, the authors used AI-based tools for grammatical refinement and stylistic polishing to improve the clarity of the presentation. Following this process, the authors conducted a thorough manual review and edit of the content to ensure its technical accuracy and original intent. The authors take full responsibility for the final content of the paper.